OpenStudy (anonymous):

Ho do i find dy/dx if y=e^x sin 2x

OpenStudy (anonymous):

taking loge both sides we get, ->lny=xsin2x differentiating both sides with respect to x we get, -> \[\frac{ 1 }{ y }\frac{ dy }{dx }=2xcos2x + \sin2x\]\[\frac{ dy }{ dx }= y (2xcos2x +\sin2x)\]\[\frac{ dy }{ dx }= e ^{xsin2x}(2xcos2x +\sin2x)\] please comment if some step is wrong